OpenStudy (anonymous):

Which fraction is the reciprocal of 4 1/9 9/37 36/1 9/4 37/9

OpenStudy (anonymous):

\[4\frac{ 1 }{ 9 }\] is a "mixed" fraction.... and can be simplified to \[\frac{ 37 }{ 9 }\] ... the "reciprocal" means its inverse, which is, \[\frac{ 9 }{ 37 }\]

Parth (parthkohli):

Rule #1: The reciprocal of \(n\) is \(\dfrac{1}{n}\). You guessed it right: the reciprocal of \(4\dfrac{1}{9}\) is \(\dfrac{1}{4\dfrac{4}{9}}\). You can simplify that further.

OpenStudy (anonymous):

thank u both Ⓐ ParthKohli & ekaknr

Parth (parthkohli):

No problem. :-)